C3H8 + 5 O2 - - - - - - - - > 3 CO2 + 4 H2O
Calculate how many molecules of CO2 are produced when 2.3 x 106 atoms of O2 are reacted.
1
Expert's answer
2020-06-26T01:04:54-0400

A mole of oxygen atoms (with the formula O) contains 6.02×1023 O atoms. A mole of oxygen molecules (formula O2) contains 6.02×1023 O2 molecules – that is, 2×6.02×1023 O atoms.

In this problem there are (2.3×106) /2 = 1.15×106 molecules of O2.

Proportion according to the reaction:

5 – 3

1.15 ×106 – x

x = 0.69 ×106 molecules of CO2

Answer: 6.9 ×105 molecules